OpenStudy (anonymous):

Ace car rental advertises that a rental car costs $25 per day plus a charge of 10 cents per mile. For the same car, Better Car rental advertises a price of $40 per day plus 5 cents per mile. What number of mile is the cost of renting a car the same at both companies ?

OpenStudy (mathstudent55):

Let's say you drive x miles. For Ace car rental, how much will the x miles cost you? For better car rental, how much will the x miles cost you?

OpenStudy (anonymous):

Ace car rental 10 cents per mile and Better car rental 5 cents per mile.

OpenStudy (mathstudent55):

Right. Now if you drive x miles, how much will that cost?

OpenStudy (anonymous):

It's asking what number of miles is the cost of renting a car the same at both companies.

OpenStudy (mathstudent55):

This is what I mean For Ace car rental: 1 mile costs 0.10 * 1 2 miles cost 0.10 * 2 x miles cost 0.01 * x = 0.10x = 0.1x For Better car rental: 1 mile costs 0.05 * 1 2 miles cost 0.05 * 2 x miles cost 0.05 * x = 0.05x

OpenStudy (mathstudent55):

Now that we know how much driving x miles will cost with each car rental company, we need to look at how much the total cost of each rental company will be. Ace charges $25 per day. That means one day plus x miles would cost 25 + 0.1x. Better charges $40 per day That means one day plus x miles would cost 40 + 0.05x. Do you follow so far?

OpenStudy (anonymous):

yes - but I don't know how to come up with how may miles for each company would be the same.

OpenStudy (mathstudent55):

We're getting there.

OpenStudy (mathstudent55):

Now notice the question we are asked. "What number of miles is the cost of renting a car the same at both companies?"

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

Since we know it costs 25 + 0.1x at Ace and 40 + 0.05x at Better, and we want the costs to be the same, we set the costs equal to each other. 25 + 0.1x = 40 + 0.05x Now we have an equation with only one unknown which we can solve. Also, notice the unknown is x which from the beginning we let it be the number of miles.

OpenStudy (mathstudent55):

Now we solve this equation. 25 + 0.1x = 40 + 0.05x

OpenStudy (mathstudent55):

Let's subtract 0.05x from both sides: 25 + 0.05x = 40 Now let's subtract 25 from both sides: 0.05x = 15 Now we divide both sides by 0.05 x = 300 The answer is 300 miles.

OpenStudy (anonymous):

Gosh !! Thanks so much !!!!!

OpenStudy (mathstudent55):

Now we check the answer. Ace: 25 + 0.1x = 25 + 0.1(300) = 25 + 30 = 55 ($55 for 1 day and 300 miles). Better: 40 + 0.05x = 40 + 0.05(300) = 40 + 15 = 55 ($55 for 1 day and 300 miles). Our answer is correct. When driving 300 miles, both rental places cost $55 for 1 day. Remember, the answer is 300 miles, not $55, since we were asked how many miles would it take to cost the same. We were not asked what the actual cost would be.

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

How about this one - If y =(x+3)2, then (-2x-6)2 would equal what?